A survey found that 60% of American victims of health care fraud are senior citizens. If 10 victims are selected at random, find

(A) the mean number of victims who are senior citizens. (

C) the probability that at most 6 victims are senior citizens.

(D) the probability that all but one victim are senior citizens.

D is the one I need most help in, I have gotten the answer "0.0016" "0.0403" and "0.999895" im very confused with it.

also can you point out what the failure and success of this problem would be (in words)
Answer #1

(n=10, p= 0.60) 10-2 Þ= 60% = 0.60, n=10. X- Binonial The pmf = P(x-x) = C) (0.60) * (0.40 (a) Mean = Elx) = np = 1 (b) PCX c